A bullet of mass 0.0018 kg moving at 472 m/s embeds itself in a large fixed piece of wood and travels 0.61 m before coming to rest. Assume that the deceleration of the bullet is constant.

1. What force is exerted by the wood on the bullet?

F =
Answer #1

Just apply kinematic equations

U = 472 m/s, V =0, s=0.61 solving we get a = 182609 m/s^2

Force exerted = -ma = 0.0018*182609 =328.69 N

Force exerted is inertia force of bullet resisted by body

We can even use impulse momentum relation.